OpenStudy (anonymous):

If x+y =1 and x^2+y^2=2, what is x/y+y/x?

OpenStudy (lgbasallote):

if x+ y = 1 then that means x = 1 - y substitute that x^2 + y^2 = 2 (1-y)^2 + y^2 = 2 1 - 2y + y^2 + y^2 = 2 1 - 2y + 2y^2 = 2 -2 + 1 - 2y + 2y^2 = 0 2y^2 - 2y - 1 = 0 you have to use quadratic formula to find y now

mathslover (mathslover):